OpenStudy (sgayton27):

Use the unit circle to find all values of θ between 0 and 2π for the following; tanθ = (-sqrt3)/3

OpenStudy (sgayton27):

\[\tan \theta=-\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (mertsj):

In which quadrants is the tangent negative?

OpenStudy (sgayton27):

tangent is negative in quadrant 2 and four

OpenStudy (mertsj):

What angle has sqrt3/3 for its tangent?

OpenStudy (sgayton27):

30 degrees

OpenStudy (mertsj):

What angle is in quadrant 2 that has a 30 degree reference angle?

OpenStudy (sgayton27):

150

OpenStudy (mertsj):

What angle in quadrant 4 has a 30 degree reference angle?

OpenStudy (sgayton27):

330 Oh i see I understand now! Thank you!

OpenStudy (mertsj):

yw